OpenStudy (anonymous):

Find the number of sides of a regular polygon if one interior angle is 60 degrees.

OpenStudy (anonymous):

use this formula k = 180(n-2)/n plug in k = 60 to get 60 = 180(n-2)/n 60n = 180(n-2) 60n = 180n - 360

OpenStudy (anonymous):

That gives you ((n - 2)180) / n, but this equals 60 in your problem, so you get ((n - 2)180)/n = 60

OpenStudy (tkhunny):

Exterior Angles ALWAYS add to 360º. Interior Angles and Exterior Angles are Supplementary. If the Interior Angle is 60º, then the Exterior Angle is 180º - 60º = 120º. If the Exterior Angle is 120º, then 360º/120º = 3 The total number of sides.

OpenStudy (anonymous):

n = 3